OpenStudy (anonymous):

HELP!!!!!A particle is dropped from a height,h above ground! If it covers 1/3rd of its distance in its last sec,find the time of flight!!!!!!HELP

OpenStudy (anonymous):

@thomaster @Kainui @karatechopper @jhonyy9 Help Guys!!

OpenStudy (kainui):

So what things do you know? The acceleration due to gravity and several equations relating position, velocity, acceleration, and time. Draw a picture and label it and try to use the equations you have along with some algebra to solve for what you need.

OpenStudy (karatechopper):

It goes 1/3 the way in a second? is that what it means?

OpenStudy (karatechopper):

and then u find the total flight?

OpenStudy (anonymous):

Thats the question.....1/3rd of its total distance in the last second!!

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I have done it in a way and found an answer.Dunno if its r8!!

OpenStudy (anonymous):

Shall I post it?

OpenStudy (anonymous):

I need another way to find the answer,easier if any!!Dunno if this's r8!!

OpenStudy (anonymous):

Any Idea @karatechopper ??

OpenStudy (karatechopper):

sorry I don't quite know.. :/

OpenStudy (anonymous):

K...tnx anyways!!

OpenStudy (anonymous):

so, we can write the equation for the last second: 0=1/3 -v1 - g/2 where i plugged 1

OpenStudy (anonymous):

plugged t =1

OpenStudy (anonymous):

now we can find v1 (the velocity 1 second before the particle hits the ground)

OpenStudy (anonymous):

v1 = 1/3 - g/2 (it will be positive since i took a negative sign)

OpenStudy (anonymous):

Why take t=1??

OpenStudy (anonymous):

i just want to find the initial velocity before the last second

OpenStudy (anonymous):

what is g in those units ?

OpenStudy (anonymous):

32.17?

OpenStudy (anonymous):

9.8

OpenStudy (anonymous):

but you gave data in yards ? :O

OpenStudy (anonymous):

M not gtin ya

OpenStudy (anonymous):

oh sorryyy grr

OpenStudy (anonymous):

No ....There's no hint of units of distance!!

OpenStudy (anonymous):

sorry sorry

OpenStudy (anonymous):

Np...cntu!!

OpenStudy (anonymous):

Did I do it in the correct way??.Hav posted my work above!!

OpenStudy (anonymous):

i might slipped somewhere though

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

anyway im not focused but this is how i would do it.. im sure there are mistakes here

OpenStudy (anonymous):

NP coolsector

OpenStudy (anonymous):

Tnx for Ur Valiant effort!!

OpenStudy (anonymous):

tnx @Coolsector ....Guess no one can help me with this!!

OpenStudy (anonymous):

tnx.... @karatechopper !!

OpenStudy (anonymous):

tnx.......@Kanui

OpenStudy (anonymous):

tnx.......@Kainui

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@Coolsector ...Okay if I close this post??

OpenStudy (anonymous):

ok so using v^2 = v0^2 + 2g(y-y0) we find the velocity before the last part of flight (that h/3) : v^2 = 2g(h/3 - h) v^2 = -4hg/3 v = sqrt(-4hg/3) also v/g=t now the last part: (1 second so t=1) 0 = h/3 + v + g/2 0 = h/3 + sqrt(-4hg/3) + g/2 h=3sqrt(6)g -15g/2

OpenStudy (anonymous):

ok i dont mind i keep on doing mistakes anyway lol

OpenStudy (anonymous):

but this is the way to the solution anyway for sure

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think I'll let it remain open for a while....Sumone's out there who will know the answer!!

OpenStudy (anonymous):

@Coolsector

OpenStudy (theeric):

My thanks to @Coolsector for the derivation of the velocity \(\frac{2}{3}\) down! \(v^2=v_0^2+2\ a\ \Delta d\) \(=v_{top}^2+2\ g\ \left(\frac{2}{3}h\right)=2\ g\ \left(\frac{2}{3}h\right)\) Now tow find the entire height: \(d=d_0+v_{\tfrac{2}{3}}\ t+\frac{1}{2}\ a\ t^2\) \(\qquad\Downarrow\) \(h=\frac{2}{3}h+v_{\tfrac{2}{3}}\ t+\frac{1}{2}\ g\ t^2\) Now you can solve for \(h\). That's what I've gotten to so far. Do you have any questions @cambrige ?

OpenStudy (theeric):

\(d=v_i\ \Delta t+\frac{1}{2}\ a\ \Delta t^2\) \(=h=(0)\Delta t+\frac{1}{2}\ g\ \Delta t^2\) So, now you can solve for the time interval for the entire fall, \(\Delta t\).

OpenStudy (foolaroundmath):

Let the total height be \(h\) and the total time of flight be \(t\). Using \(\displaystyle d = d_{0}+v_{0}t +\frac{1}{2}at^2\) for the entire motion, \(d_0 = 0\) and \(v_0 = 0\), So \(\displaystyle h=\frac{1}{2}gt^2\) ...... (1) Now consider the motion for the time \(t-1\) and use the same equation. Here \(d = 2h/3\) \(\displaystyle \frac{2h}{3} = 0 + 0\cdot (t-1) + \frac{1}{2}g(t-1)^2\) Substitute the value of \(h\) obtained in (1) \(\displaystyle \frac{1}{3}\cancel{g}t^2 = \frac{1}{2}\cancel{g}(t-1)^2\) Now take squareroot on both sides \(\displaystyle \frac{\pm t}{\sqrt 3} = \frac{t-1}{\sqrt 2}\) or \(\displaystyle t = 3 \pm \sqrt 6\) Since the time of flight has to be atleast greater than 1 sec (since it covers h/3 in the last second), so the only answer possible is \(t = 3+\sqrt 6\)

OpenStudy (anonymous):

@FoolAroundMath :)

OpenStudy (anonymous):

So My ans is right,I guess...

OpenStudy (anonymous):

No @theEric....No more Q's....Tnx for the help!!

OpenStudy (anonymous):

Tnx to @FoolAroundMath